OpenStudy (yttrium):

The greatest integer function,[x] , is equal to the largest integer that is smaller than or equal to x. In the domain [0,100], how many solutions are there to x^2 - [x] x = 88.25?

OpenStudy (anonymous):

Similar problem here: http://openstudy.com/study#/updates/53081551e4b00fb1e23842b8

OpenStudy (anonymous):

Basically, recall that: \[ 0\leq x-\lfloor x \rfloor <1 \]

OpenStudy (anonymous):

\[ 0\leq \frac{88.25}{x} < 1 \]

OpenStudy (anonymous):

This means we at least can realize:\[ 88.25<x\leq 100 \]It's not feasible otherwise.

OpenStudy (anonymous):

If we partition the interval \((88.25,100]\) based on \(\lfloor x \rfloor \), we can find a single solution in each partion.

OpenStudy (anonymous):

Except maybe \(\lfloor x\rfloor = 100\) and \(\lfloor x\rfloor =88\). Those we'd have to double check.

OpenStudy (anonymous):

11.75 is your answer, idk how but believe in the bat.

OpenStudy (anonymous):

How can you have 0.75 of a solution? lol

OpenStudy (anonymous):

12

OpenStudy (anonymous):

okay hold on a second though.

OpenStudy (anonymous):

I agree that \(\lfloor x \rfloor =89, 90, 91, \dots 99\) will have solutions for sure.

OpenStudy (anonymous):

So we have 11 solutions. But then we need to figure out if there is a solution in \(\lfloor x\rfloor = 88\).

OpenStudy (anonymous):

\[ x(x-88) = 88.25 \]

OpenStudy (anonymous):

Hmm, that gives \(x = 88.9917\).... so I agree there is a 12th solution.

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

okay now I'm wondering...

OpenStudy (anonymous):

If we let \(0\leq b < 1\) and have \(n\)... So instead of \(88.25\) we'd say \(b=0.25\), \(n=88\), \(88.25 = n+b\)

OpenStudy (anonymous):

\[ x(x-n) = n+b \] will this always have a solution?

OpenStudy (anonymous):

That is, a real solution \(x_0\) that we know is \(n+b<x_0<n+1\)

OpenStudy (anonymous):

let x=n+d, where n is an integer and 0<d<1 then x^2-x[x] = x(x-[x]) = (n+d)(n+d-n) = (n+d)d = 88.25 which will have solutions from n=88 to n=100 thus the number of solutions is 13

OpenStudy (anonymous):

for n=88, we have (88+d)d = 88.25 88d+d^2 = 88.25 let f(d)=88d+d^2 f(0)=0 f(1)=89 since f is continuous, for some value d between 0 and 1, f(d)=88.25

OpenStudy (anonymous):

Well, \(d\) can't be under \(0.25\)